OpenStudy (anonymous):

Find or evaluate the integral. (Complete the square, if necessary. Remember to use ln |u| where appropriate.) (8x − 2)/(x^2 + 2x + 2) dx

OpenStudy (anonymous):

hey can you help me on this one too

OpenStudy (anonymous):

yeah I'll try

OpenStudy (anonymous):

okiii

myininaya (myininaya):

let me know if you need more help

OpenStudy (zarkon):

you have a 'typo'

OpenStudy (anonymous):

hmm um at the end is that thue answer...

OpenStudy (zarkon):

\[x^2 + 2x + 2=x^2+2x+1-1+2\] \[=(x^2+2x+1)+1=(x+1)^2+1\]

myininaya (myininaya):

you are right i added 1 twice

OpenStudy (anonymous):

(x+1)^2+3 is what myininaya has

OpenStudy (zarkon):

I would break it up into two integrals

myininaya (myininaya):

yes i added 1 twice its a mistake it should be what zarkon has

myininaya (myininaya):

so we should \[2\int\limits_{}^{}\frac{4\tan(\theta)-5}{\sec^2(\theta)} d \theta\]

OpenStudy (anonymous):

Yeah Zarkon that's what I was thinking too... \[\int\limits_{}^{} \frac{4(2x+2)}{x^2+2x+2} \ + \int\limits_{}^{ }\frac{-10}{x^2 + 2x+2} \]

OpenStudy (zarkon):

that works nice

myininaya (myininaya):

\[2(\int\limits_{}^{}\frac{4\tan(\theta)}{\sec^2(\theta)} d \theta-\int\limits_{}^{}\frac{5}{\sec^2(\theta)} d \theta)\]

myininaya (myininaya):

\[\int\limits_{}^{}2\sin(\theta)\cos(\theta) d \theta-10\int\limits_{}^{}\cos^2(\theta) d \theta\] \[\int\limits_{}^{}\sin(2\theta) dt heta-10 \int\limits_{}^{}\frac{1}{2}(1+\cos(2\theta))d \theta\]

OpenStudy (anonymous):

\[let u = x^2 + 2z + 2; du = (2x + 2)dx \] So for the left integral we have: \[4\int\limits_{}^{}\ \frac{du}{u} \ = 4\ln \left| u \right|\]

myininaya (myininaya):

\[=-\frac{1}{2} \cos(2 \theta)-5(\theta+\frac{1}{2}\sin( 2 \theta))+C\]

myininaya (myininaya):

now just put it back in terms of x

OpenStudy (anonymous):

Right: \[4\ln \left| x^2 + 2x +2 \right|\]

OpenStudy (anonymous):

For the right integral we have to complete the square in the denominator: \[x^2 + 2x + 2 = (x + 1)^{2} + 1\]

OpenStudy (anonymous):

so the final answer is what myininaya wrote out but just put x in place of the theta symbol?

OpenStudy (zarkon):

not exactly..it is a little more complicated than that

OpenStudy (anonymous):

Therefore the integral will be:\[-10\int\limits_{}^{}\ \frac{dx}{(x+1)^{2}+1} \]

myininaya (myininaya):

x does not equal theta

myininaya (myininaya):

\[\tan(\theta)=x+1\] thats x and theta's relationship

myininaya (myininaya):

\[\theta=\tan^{-1}(x+1)\]

OpenStudy (anonymous):

Now we have to figure out a way to utilize arctan(x) derivative so we must set an intermediate variable like a = x + 1 and if we take the derivative da = dx. So our integral becomes

myininaya (myininaya):

\[\tan(\theta)=\frac{opposite}{adjacent}=\frac{x+1}{1}=> hyp=\sqrt{(x+1)^2+1}\] \[\cos(\theta)=\frac{1}{\sqrt{(x+1)^2+1}}, \sin(\theta)=\frac{x+1}{\sqrt{(x+1)^2+1}}\]

OpenStudy (anonymous):

\[-10\int\limits_{}^{}\ \frac{da}{a^2 + 1} \ - -10 \tan^{-1} a= -10\tan^{-1} (x+1)\]

OpenStudy (anonymous):

Therefore the final answer will be:\[4\ln \left| x^2 + 2x+2 \right| - 10\tan^{-1} (x+1)\]

myininaya (myininaya):

\[=-\frac{1}{2} \cos(2 \theta)-5(\theta+\frac{1}{2}\sin( 2 \theta))+C \] \[=\frac{-1}{2}(\cos^2(\theta)-\sin^2(\theta))-5 \theta-\frac{5}{2}2\sin(\theta)\cos(\theta)+C\] \[=\frac{-1}{2}(\frac{1}{(x+1)^2+1}-\frac{(x+1)^2}{(x+1)^2+1})-5\tan^{-1}(x+1)-5\frac{x+1}{(x+1)^2+1}+C\]

myininaya (myininaya):

and that can be cleaned up

OpenStudy (anonymous):

wait so whos rightis nayeaddo right or you myinaya

myininaya (myininaya):

go through and check our work and you can come to that conclusion or you can enter it into wolfram to see who;s answer is right

myininaya (myininaya):

honestly after that i don't feel like checking my work sorry

OpenStudy (zarkon):

lol

OpenStudy (zarkon):

I'm telling you...your arms are going to fall off

OpenStudy (anonymous):

Yeah myininaya your method is very inricate lol...I mean it involves a lot

myininaya (myininaya):

based on wolfram nayeaddo has it

myininaya (myininaya):

use that site to check your work

myininaya (myininaya):

its awesome

myininaya (myininaya):

nayeddo i learned a new word thanks

OpenStudy (anonymous):

cool what word?

myininaya (myininaya):

inricate thats weird it says i spelled it wrong

OpenStudy (anonymous):

yeah there is at before the r

OpenStudy (anonymous):

and nice!

myininaya (myininaya):

i became your fan naye because you won against me this round but i will win next time

OpenStudy (anonymous):

hahaha...I like that competition lol I also became your fan so I guess its on..but like I said your methods are interesting too so we'll see

myininaya (myininaya):

i guess it would had been better if i separated the fraction at the beginning

myininaya (myininaya):

or easier

OpenStudy (anonymous):

yeah i think especially for the one who asked the question

OpenStudy (anonymous):

But it also makes the math easier to understand.. less complicated

myininaya (myininaya):

but complications make things more dreamy

myininaya (myininaya):

that's not suppose to make too much sense so don't read too much in that statement

OpenStudy (anonymous):

yeah...I kind of know what you mean though